OpenStudy (anonymous):

help =[ use factoring to solve the quadratic equation:9x^2=30x-25 ^means exponent/power

OpenStudy (anonymous):

9x^2=30x-25 9x^2-30x+25=0 9x^2-15x-15x+25=0 3x(3x-5)-5(3x-5)=0 (3x-5)(3x-5)=0